### v1

# On Touchard’s Identity: Generalizations and Related Results

Abstract
Starting with a known polynomial identity, we derive two generalizations of Touchard’s identity concerning Catalan numbers; one obtained using the Beta function and the other via a connection with Stirling numbers of the second kind. We subsequently establish several new combinatorial identities.
2020 Mathematics Subject Classification: Primary 05A19; Secondary 05A10.
Keywords: Catalan number, Touchard’s identity, harmonic number, Stirling number of the second kind, binomial transform.
1 Introduction
Let $C_{k}=\binom{2k}{k}/(k+1)$ be the $k$th Catalan number. The celebrated Touchard’s identity [4, 5, 6, 7]:
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{n}{{2k}}2^{n-2k}C_{k}}=C_{n+1},$ (1)
is subsumed at $x=0$ by the following identity [2, Equation (54)]:
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}\left({1-x}\right)^{2k}}=\sum_{k=0}^{n}{\binom{{n}}{k}2^{-k}C_{k+1}\left({1-x}\right)^{k}x^{n-k}}.$ (2)
Based on (2), we will derive the following generalization of (1):
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{n-2k}C_{k}\binom{{2k+r}}{s}^{-1}}=\frac{s}{{n+r}}\sum_{k=0}^{n}{\binom{{n}}{k}2^{n-k}C_{k+1}\binom{{n+r-1}}{{k+r-s}}^{-1}},$ (3)
which holds for complex numbers $r$ and $s$ that are not negative integers.
Note that (3) reduces to (1) when $s=0$, since
$\lim_{s\to 0}s\binom{p-1}{p-s}^{-1}=p.$ (4)
By linking (2) with Stirling numbers of the second kind, defined for non-negative integers $m$ and $n$ through the generating function
[…]
and having the explicit representation
[…]
we will derive another generalization of Touchard’s identity, namely,
[…]
which holds for every non-negative integer $r$.
Identity (2) has further interesting immediate consequences. Evaluation at $x=2$ produces the self-inverse sequence binomial transform identity
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-2k}C_{k+1}}=2^{-2n}C_{n+1},$ (8)
which was reported by Donaghey [3]. Setting x=1x=-1 and $x=3$, in turn, in (2), produces the following interesting identity:
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}C_{k+1}}=(-3)^{n}\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}3^{-k}C_{k+1}}=(-1)^{n}\sum_{k=0}^{n}{\binom{{n}}{{2k}}C_{k}}.$ (9)
Binomial coefficients are defined, for non-negative integers $i$ and $j$, by
$\binom{i}{j}=\begin{cases}\dfrac{{i!}}{{j!(i-j)!}},&\text{$i\geq j$};\\ 0,&\text{$i<j$}.\end{cases}$ (10)
The definition is extended to complex numbers $r$ and $s$ by
$\binom{r}{s}=\frac{{\Gamma(r+1)}}{{\Gamma(s+1)\Gamma(r-s+1)}},$ (11)
where the Gamma function, $\Gamma(z)$, is defined for $\Re(z)>0$ by
$\Gamma(z)=\int_{0}^{\infty}{e^{-t}t^{z-1}dt}=\int_{0}^{\infty}{\left({\log(1/t)}\right)^{z-1}dt},$ (12)
and is extended to the rest of the complex plane, excluding the non-positive integers, by analytic continuation.
Relationships between binomial coefficients can be established by using (11), together with the following identities:
$\Gamma\left({z+\frac{1}{2}}\right)=\sqrt{\pi}\,2^{-2z}\binom{{2z}}{z}\,\Gamma\left({z+1}\right),$ (13)
and
$\Gamma\left({-z+\frac{1}{2}}\right)=(-1)^{z}\,2^{2z}\,\binom{{2z}}{z}^{-1}\frac{{\sqrt{\pi}}}{{\Gamma\left({z+1}\right)}}.$ (14)
Harmonic numbers, $H_{n}$, and odd harmonic numbers, $O_{n}$, are defined for non-negative integers $n$ by
$H_{n}=\sum_{k=1}^{n}\frac{1}{k},\quad O_{n}=\sum_{k=1}^{n}\frac{1}{2k-1},\quad H_{0}=0=O_{0}.$ (15)
We will make frequent use of the following identity
$H_{n-1/2}=2O_{n}-2\ln 2,$ (16)
extending harmonic numbers to half-integer arguments and which is a consequence of the link between harmonic numbers, odd harmonic numbers and the digamma function.
2 A generalization of Touchard’s identity and related results
Our first generalization of Touchard’s identity (1) is stated in Theorem 1. We require the integration formulas stated in Lemma 1, which are variations on the well-known Beta function.
Lemma 1.
If $r$ and $s$ are complex numbers, then
$\int_{0}^{1}{y^{r+k-s}\left({1-y}\right)^{s-1}dy}=\frac{1}{s}\binom{{k+r}}{s}^{-1}$ (17)
and
$\int_{0}^{1}{y^{s+k}\left({1-y}\right)^{r-k}dy}=\frac{1}{{r+s+1}}\binom{{r+s}}{{k+s}}^{-1}.$ (18)
Theorem 1.
If $r$ and $s$ are complex numbers that are not negative integers, then
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}\binom{{2k+r}}{s}^{-1}}=\frac{s}{{n+r}}\sum_{k=0}^{n}{\binom{{n}}{k}2^{-k}C_{k+1}\binom{{n+r-1}}{{k+r-s}}^{-1}}.$ (19)
Proof.
Using (2), write
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}y^{r-s+2k}\left({1-y}\right)^{s-1}}=\sum_{k=0}^{n}{\binom{{n}}{k}2^{-k}C_{k+1}y^{k+r-s}\left({1-y}\right)^{n-k+s-1}},$ (20)
and integrate term-wise from 0 to 11, using Lemma 1. ∎
The rest of this section is devoted to deriving some consequences of (3).
Proposition 1.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-6k}C_{k}C_{2k}}=\frac{1}{{\left({n+1}\right)2^{2n}}}\sum_{k=0}^{n}{\left({2\left({n-k}\right)+1}\right)\binom{{2\left({n-k}\right)}}{{n-k}}\binom{{2k}}{k}2^{-k}C_{k+1}}$ (21)
and
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{2k}\frac{{C_{k}}}{{C_{2k}}}}=3\binom{{2n}}{n}^{-1}\sum_{k=0}^{n}{\frac{{2^{k}\left({k+1}\right)}}{{\left({2\left(n-k\right)-1}\right)\left({2\left(n-k\right)-3}\right)}}\binom{{2\left({n-k}\right)}}{{n-k}}C_{k+1}}.$ (22)
Proof.
To derive (21), set $r=1$ and $s=3/2$ in (3) and use
$\binom{{2k+1}}{{3/2}}=\frac{4}{{3\sqrt{\pi}}}\frac{{\Gamma\left({2k+2}\right)}}{{\Gamma\left({2k+1/2}\right)}}=\frac{4}{{3\pi}}\frac{{2^{4k}}}{{C_{2k}}}$ (23)
and
$\binom{{n}}{{k-1/2}}=\frac{{\Gamma\left({n+1}\right)}}{{\Gamma\left({k+\frac{1}{2}}\right)\Gamma\left({n-k+\frac{3}{2}}\right)}}=\frac{{2^{2n+1}}}{{\pi\left({2\left({n-k}\right)+1}\right)}}\binom{{n}}{k}\binom{{2\left({n-k}\right)}}{{n-k}}^{-1}\binom{{2k}}{k}^{-1}.$ (24)
Identity (22) is obtained by setting r=1/2r=-1/2 and s=3/2s=-3/2 in (3) and using
$\binom{{2k-1/2}}{{-3/2}}=-2^{-4k-1}C_{2k}$ (25)
and
$\binom{{n-3/2}}{{k+1}}=\frac{{2^{-2k}}}{2}\frac{{\left({2\left({n-k}\right)-1}\right)\left({2\left({n-k}\right)-3}\right)}}{{\left({2n-1}\right)\left({k+1}\right)}}\binom{{2n}}{n}\binom{{n}}{k}\binom{{2\left({n-k}\right)}}{{n-k}}$. (26)
Proposition 2.
If $n$ is a non-negative integer and $r$ is a complex number that is not a non-positive integer, then
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}\frac{{2^{-2k}}}{{2k+r}}\,C_{k}}=\frac{{1}}{{n+1}}\binom{{n+r}}{{n+1}}^{-1}\sum_{k=0}^{n}{2^{-k}C_{k+1}\binom{{k+r-1}}{k}}$. (27)
In particular,
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}\frac{{2^{-2k}}}{{2k+1}}}\,C_{k}=\frac{1}{{n+1}}\sum_{k=0}^{n}{2^{-k}C_{k+1}}$ (28)
and
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}\frac{{2^{-2k}}}{{4k+1}}}\,C_{k}=\frac{{2^{2n}}}{{2n+1}}\binom{{2n}}{n}^{-1}\sum_{k=0}^{n}{2^{-3k}\binom{{2k}}{k}C_{k+1}}.$ (29)
Proof.
Set $s=1$ in (3) and use
$\binom{{n+r-1}}{{k+r-1}}\binom{{k+r-1}}{k}=\binom{{n}}{k}\binom{{n+r-1}}{n}$. (30)
Identity (28) corresponds to setting $r=1$ in (27) while Identity (29) is obtained by evaluating (27) at $r=1/2$ and making use of
$\binom{{n+1/2}}{{n+1}}=\frac{{2n+1}}{{2^{2n+1}}}C_{n}$ (31)
and
$\binom{{k-1/2}}{k}=2^{-2k}\binom{{2k}}{k}$. (32)
Proposition 3.
If $n$ is a non-negative integer and $s$ is a complex number that is not a negative integer, then
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}}2^{-2k}C_{k}\binom{{2k+s}}{s}^{-1}=\frac{s}{{n+s}}\sum_{k=0}^{n}{\binom{{n}}{k}2^{-k}C_{k+1}\binom{{n-1+s}}{k}^{-1}}.$ (33)
In particular,
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}}2^{2k}C_{k}\binom{{4k}}{{2k}}^{-1}$ (34)
$\qquad=\frac{2^{n}3n}{{\left({n+1}\right)\left({n+2}\right)}}-\sum_{k=0}^{n-2}{\frac{{2^{k}}}{{2n-2k-1}}\binom{{n}}{k}^{2}C_{k+1}\binom{2k}{k}^{-1}\binom{2n}{2k}^{-1}}.$ (35)
Proof.
Set $r=s$ in (3). To derive (35), set s=1/2s=-1/2 in (33), write
$\sum_{k=0}^{n}{\binom{{n}}{k}2^{-k}C_{k+1}\binom{{n-2+1/2}}{k}^{-1}}$ $=\sum_{k=0}^{n-2}{\binom{{n}}{k}2^{-k}C_{k+1}\binom{{n-2+1/2}}{k}^{-1}}$ (36)
$\qquad+\binom{{n}}{{n-1}}2^{-(n-1)}C_{n}\binom{{n-2+1/2}}{{n-1}}^{-1}$ (37)
$\qquad\qquad+2^{-n}C_{n+1}\binom{{n-2+1/2}}{n}^{-1}$ (38)
and use
$\binom{{n-2+1/2}}{k}=2^{-2k}\binom{{2n-3}}{{2k}}\binom{{2k}}{k}\binom{{n-2}}{k}^{-1}$ (39)
and
$\binom{{2k-1/2}}{{-1/2}}=\binom{{2k-1/2}}{{2k}}=2^{-4k}\binom{{4k}}{{2k}}$. (40)
Proposition 4.
If $n$ and $r$ are non-negative integers and $s$ is a complex number such that sns-n is not a negative integer, then
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}\binom{{n-s-1}}{{2k+r}}^{-1}}=(-1)^{r-1}\,\frac{{n-s}}{{r+s}}\,\sum_{k=0}^{n}{\binom{{n}}{k}2^{-k}C_{k+1}\binom{{r+s-1}}{{k+r}}^{-1}}.$ (41)
In particular,
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\frac{{2^{2k}}}{{\left({2n-4k-1}\right)\left({2n-4k-3}\right)}}\binom{{n}}{{2k}}^{2}\binom{{2k}}{k}\binom{{2n}}{{4k}}^{-1}\frac{1}{{C_{2k}}}}=(-1)^{n}\,\frac{{2n+1}}{3}.$ (42)
Proof.
Write rnr-n for $r$ and sns-n for $s$ in (3) and write $r+s$ for $r$ in the resulting equation, noting that
$\binom{2k+r+s-n}{2k+r}=(-1)^{r}\binom{n-s-1}{2k+r}$. (43)
To derive (42), set $r=2$ and s=1/2s=-1/2 in (41) and use
$\binom{{1/2}}{{k+2}}=\frac{{(-1)^{k+1}}}{{2^{2k+3}}}C_{k+1}$ (44)
and
$\binom{{n-1/2}}{{2k+2}}=\frac{{2^{-4k}}}{8}\frac{{\left({2n-4k-1}\right)\left({2n-4k-3}\right)}}{{k+1}}\binom{{2n}}{{4k}}\binom{{n}}{{2k}}^{-1}C_{2k}.$ (45)
Proposition 5.
If $n$ is a non-negative integer and $r$ and $s$ are complex numbers that are not negative integers, then
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}H_{2k+r-s}\binom{{2k+r}}{s}^{-1}}$ (46)
$\qquad=\frac{s}{{n+r}}\sum_{k=0}^{n}{\binom{{n}}{k}2^{-k}C_{k+1}\left({H_{k+r-s}+H_{s-1}-H_{n+s-k-1}}\right)\binom{{n+r-1}}{{k+r-s}}^{-1}}.$ (47)
In particular,
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}H_{2k+r}}=2^{-n}C_{n+1}H_{n+r}-\sum_{k=0}^{n-1}{\binom{{n}}{k}\frac{2^{-k}}{n-k}C_{k+1}\binom{{n+r}}{{k+r}}^{-1}}.$ (48)
Proof.
Differentiate (3) with respect to $s$ to obtain
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}\left({H_{2k+r-s}-H_{s}}\right)\binom{{2k+r}}{s}^{-1}}$ (49)
$\qquad=-\frac{1}{{n+r}}\sum_{k=0}^{n}{\binom{{n}}{k}2^{-k}C_{k+1}\binom{{n+r-1}}{{k+r-s}}^{-1}}$ (50)
$\qquad\qquad+\frac{s}{{n+r}}\sum_{k=0}^{n}{\binom{{n}}{k}2^{-k}C_{k+1}\left({H_{k+r-s}-H_{n+s-k-1}}\right)\binom{{n+r-1}}{{k+r-s}}^{-1}},$ (51)
from which (47) follows after using (3) again. To derive (48), note that $sH_{s-1}=sH_{s}-1$, write the right-hand side of (47) as
$\frac{1}{{n+r}}\sum_{k=0}^{n-1}{\binom{{n}}{k}2^{-k}C_{k+1}\binom{{n+r-1}}{{k+r-s}}^{-1}\left({sH_{k+r-s}+sH_{s}-1-sH_{n+s-k-1}}\right)}$ (52)
$\qquad+\frac{1}{{n+r}}2^{-n}sC_{n+1}H_{n+r-s}\binom{{n+r-1}}{{n+r-s}}^{-1}$ (53)
and take the limit as $s$ approaches zero, using (4). ∎
Proposition 6.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}H_{2k}}=2^{-n}C_{n+1}H_{n}-\sum_{k=0}^{n-1}{\frac{{2^{-k}C_{k+1}}}{{n-k}}}$ (54)
and
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}O_{2k}}$ (55)
$\qquad=2^{-n}C_{n+1}O_{n}-2^{-n-1}\sum_{k=1}^{n}{\frac{{2^{3k}}}{k}\binom{{n}}{k}^{2}\binom{{2k}}{k}^{-1}\binom{{2n}}{{2k}}^{-1}C_{n-k+1}}.$ (56)
Proof.
Set $r=0$ in (48) to obtain (54), and r=1/2r=-1/2 to derive (56) after some algebra. ∎
Proposition 7.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-6k}C_{k}C_{2k}O_{2k}}$ (57)
$\qquad=\frac{1}{{2^{2n}\left({n+1}\right)}}\sum_{k=0}^{n}{\left({2\left({n-k}\right)+1}\right)\binom{{2\left({n-k}\right)}}{{n-k}}\binom{{2k}}{k}2^{-k}C_{k+1}\left({O_{k}-O_{n-k+1}+1}\right)}.$ (58)
Proof.
Set $r=1$ and $s=3/2$ in (47) and equate rational parts after using (23) and (24). ∎
3 A generalization involving Stirling numbers of the second kind
Lemma 2.
If $m$ and $u$ are non-negative integers, then
[…]
Proof.
Since
$\left({1-e^{x}}\right)^{u}=\sum_{p=0}^{u}{(-1)^{p}\binom{{u}}{p}e^{px}},$ (60)
we have
$\frac{{d^{m}}}{{dx^{m}}}{\left({1-e^{x}}\right)^{u}}=\sum_{p=0}^{u}{(-1)^{p}\binom{{u}}{p}p^{m}e^{px}};$ (61)
and hence (59).
Theorem 2.
If $n$ and $r$ are non-negative integers, then
[…]
Proof.
By writing $1/x$ for $x$, Identity (2) can also be written as
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}x^{n-2k}}=\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}2^{-k}C_{k+1}\left({1-x}\right)^{n-k}}.$ (63)
Now replace $x$ with $\exp x$, differentiate the resulting expression $r$ times with respect to $x$ and evaluate at $x=0$, making use of Lemma 2. ∎
4 Related results via binomial transform
Two sequences of complex numbers $(s_{k})$ and $(\sigma_{k})$, $k=0,1,2,\ldots$, are called a binomial transform pair if
$s_{n}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\sigma_{k}},$ (64)
for every non-negative integer $n$. It is known that
$s_{n}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\sigma_{k}}\iff\sigma_{n}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}s_{k}}.$ (65)
Lemma 3 ([1, Lemma 2.1]).
Let $\left\{(a_{k}),(\alpha_{k})\right\}$, $k=0,1,2,\ldots$, be a binomial-transform pair. Let $\mathcal{L}_{x}$ be a linear operator defined by $\mathcal{L}_{x}(x^{j})=a_{j}$ for every complex number $x$ and every non-negative integer $j$. Then $\mathcal{L}_{x}((1-x)^{j})=\alpha_{j}$.
Theorem 3.
Let $\left\{(a_{k}),(\alpha_{k})\right\}$, $k=0,1,2,\ldots$, be a binomial-transform pair. Then
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}2^{-k}C_{k+1}a_{n-k}}=\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}\alpha_{n-2k}}.$ (66)
Proof.
Write (63) as
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}2^{-k}C_{k+1}x^{n-k}}=\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}\left({1-x}\right)^{n-2k}}$ (67)
and apply Lemma 3. ∎
Remark 1.
It is sometimes useful to write (66) in the following form:
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{n-2k}C_{k}a_{n-2k}}$ (68)
$\qquad=\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{2k}C_{n-2k+1}\alpha_{2k}}-\sum_{k=1}^{\left\lceil{n/2}\right\rceil}{\binom{{n}}{{2k-1}}2^{2k-1}C_{n-2k+2}\alpha_{2k-1}}.$ (69)
Proposition 8.
If $n$ is a non-negative integer, then
[…]
Proof.
Use the binomial transform pair $\{a_{k}=1,\alpha_{k}=\delta_{k0}\}$ in (66). ∎
Remark 2.
Identity (70) was obtained by Suleiman and Sury [8].
Proposition 9.
If $n$ is a non-negative integer, then
[…]
and
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}\frac{{2^{n-2k}}}{{n-k+1}}\,C_{k+1}H_{n-k}}$ (72)
$\qquad=\begin{cases}-2\sum\limits_{k=0}^{(n-1)/2}{\dbinom{{n}}{{2k}}\dfrac{{2^{-2k}}}{{n-2k+1}}\,C_{k}O_{(n-2k+1)/2}},&\text{if $n$ is odd;}\\ 0,&\text{if $n$ is even.}\end{cases}$ (73)
Proof.
The second identity in [1, Proposition 13.3] is
[…]
where the binomial transform pair $(a_{k})$ and $(\alpha_{k})$, $k=0,1,2,\ldots$, where
$a_{k}=\frac{2^{k}}{k+1}\text{ and }\alpha_{k}=\frac{1+(-1)^{k}}{2\,\left(k+1\right)},$ (75)
can be identified. Use these in (66) to obtain (71). From [1, Proposition 13.4]:
[…]
we recognize the binomial transform pair $(a_{k})$ and $(\alpha_{k})$, $k=0,1,2,\ldots$, where
$a_{k}=\frac{2^{k}H_{k}}{k+1}\text{ and }\alpha_{k}=-\frac{1-(-1)^{k}}{\,\left(k+1\right)}\,O_{(k+1)/2}.$ (77)
Use of these in (66) gives (73). ∎
Proposition 10.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}\frac{{2^{2\left({n-2k}\right)}}}{{n-2k+1}}\,C_{k}}=\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}\frac{{2^{2k}}}{{2k+1}}\,C_{n-2k+1}}$ (78)
and
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}\frac{{2^{2\left({n-2k}\right)}}}{{n-2k+1}}\,C_{k}H_{n-2k}}=\frac{1}{2}\sum_{k=0}^{\left\lceil{n/2}\right\rceil}{\binom{{n}}{{2k-1}}\frac{{2^{2k}}}{k}\,C_{n-2k+2}O_{k}}.$ (79)
Proof.
Use (75) in (69) to obtain (78) and (76) to get (79). ∎