### v2

# On Touchard’s Identity: Generalizations and Related Results

Abstract
Starting with a known polynomial identity, we derive two generalizations of Touchard’s identity concerning Catalan numbers; one obtained using the Beta function and the other via a connection with Stirling numbers of the second kind. We subsequently establish several new combinatorial identities.
2020 Mathematics Subject Classification: Primary 05A19; Secondary 05A10.
Keywords: Catalan number, Touchard’s identity, Knuth’s old sum, harmonic number, Stirling number of the second kind, binomial transform.
1 Introduction
Let $C_{k}=\binom{2k}{k}/(k+1)$ be the $k$th Catalan number. The celebrated Touchard’s identity [4, 6, 7, 8]:
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{n}{{2k}}2^{n-2k}C_{k}}=C_{n+1},$ (1)
is subsumed at $x=0$ by the following identity [2, Equation (54)]:
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}\left({1-x}\right)^{2k}}=\sum_{k=0}^{n}{\binom{{n}}{k}2^{-k}C_{k+1}\left({1-x}\right)^{k}x^{n-k}}.$ (2)
Based on (2), we will derive the following generalization of (1):
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{n-2k}C_{k}\binom{{2k+r}}{s}^{-1}}=\frac{s}{{n+r}}\sum_{k=0}^{n}{\binom{{n}}{k}2^{n-k}C_{k+1}\binom{{n+r-1}}{{k+r-s}}^{-1}},$ (3)
which holds for complex numbers $r$ and $s$ that are not negative integers.
Note that (3) reduces to (1) when $s=0$, since
$\lim_{s\to 0}s\binom{p-1}{p-s}^{-1}=\lim_{s\to 0}s\,\frac{p}{s}\binom{p}{p-s}^{-1}=p.$ (4)
By linking (2) with Stirling numbers of the second kind, defined for non-negative integers $m$ and $n$ through the generating function
[…]
and having the explicit representation
[…]
we will derive another generalization of Touchard’s identity, namely,
[…]
which holds for every non-negative integer $r$.
Identity (2) has further interesting immediate consequences. Evaluation at $x=2$ produces the following self-inverse sequence binomial-transform identity:
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-2k}C_{k+1}}=2^{-2n}C_{n+1},$ (8)
which was reported by Donaghey [3]. Setting x=1x=-1, and $x=3$, in turn, in (2), produces the following interesting identity:
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}C_{k+1}}=(-3)^{n}\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}3^{-k}C_{k+1}}=(-1)^{n}\sum_{k=0}^{n}{\binom{{n}}{{2k}}C_{k}}.$ (9)
Binomial coefficients are defined, for non-negative integers $i$ and $j$, by
$\binom{i}{j}=\begin{cases}\dfrac{{i!}}{{j!(i-j)!}},&\text{$i\geq j$};\\ 0,&\text{$i<j$}.\end{cases}$ (10)
The definition is extended to complex numbers $r$ and $s$ by
$\binom{r}{s}=\frac{{\Gamma(r+1)}}{{\Gamma(s+1)\Gamma(r-s+1)}},$ (11)
where the Gamma function, $\Gamma(z)$, is defined for $\Re(z)>0$ by
$\Gamma(z)=\int_{0}^{\infty}{e^{-t}t^{z-1}dt}=\int_{0}^{\infty}{\left({\log(1/t)}\right)^{z-1}dt},$ (12)
and is extended to the rest of the complex plane, excluding the non-positive integers, by analytic continuation.
Relationships between binomial coefficients can be established by using (11), together with the following identities:
$\Gamma\left({z+\frac{1}{2}}\right)=\sqrt{\pi}\,2^{-2z}\binom{{2z}}{z}\,\Gamma\left({z+1}\right),$ (13)
and
$\Gamma\left({-z+\frac{1}{2}}\right)=(-1)^{z}\,2^{2z}\,\binom{{2z}}{z}^{-1}\frac{{\sqrt{\pi}}}{{\Gamma\left({z+1}\right)}}.$ (14)
Harmonic numbers, $H_{n}$, and odd harmonic numbers, $O_{n}$, are defined for non-negative integers $n$ by
$H_{n}=\sum_{k=1}^{n}\frac{1}{k},\quad O_{n}=\sum_{k=1}^{n}\frac{1}{2k-1},\quad H_{0}=0=O_{0}.$ (15)
We will make frequent use of the following identity:
$H_{n-1/2}=2O_{n}-2\ln 2,$ (16)
which extends harmonic numbers to half-integer arguments and which is a consequence of the link between harmonic numbers, odd harmonic numbers and the digamma function.
2 A generalization of Touchard’s identity and related results
Our first generalization of Touchard’s identity (1) is stated in Theorem 1. We require the integration formulas stated in Lemma 1, which are variations on the well-known Beta function.
Lemma 1.
If $r$ and $s$ are complex numbers, then
$\int_{0}^{1}{y^{r+k-s}\left({1-y}\right)^{s-1}dy}=\frac{1}{s}\binom{{k+r}}{s}^{-1}$ (17)
and
$\int_{0}^{1}{y^{s+k}\left({1-y}\right)^{r-k}dy}=\frac{1}{{r+s+1}}\binom{{r+s}}{{k+s}}^{-1}.$ (18)
Theorem 1.
If $r$ and $s$ are complex numbers that are not negative integers, then
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}\binom{{2k+r}}{s}^{-1}}=\frac{s}{{n+r}}\sum_{k=0}^{n}{\binom{{n}}{k}2^{-k}C_{k+1}\binom{{n+r-1}}{{k+r-s}}^{-1}}.$ (19)
Proof.
Using (2), write
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}x^{r-s+2k}\left({1-x}\right)^{s-1}}=\sum_{k=0}^{n}{\binom{{n}}{k}2^{-k}C_{k+1}x^{k+r-s}\left({1-x}\right)^{n-k+s-1}},$ (20)
and integrate term-wise from 0 to 11, using Lemma 1. ∎
The rest of this section is devoted to deriving some consequences of (3).
Proposition 1.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-6k}C_{k}C_{2k}}=\frac{1}{{\left({n+1}\right)2^{2n}}}\sum_{k=0}^{n}{\left({2\left({n-k}\right)+1}\right)\binom{{2\left({n-k}\right)}}{{n-k}}\binom{{2k}}{k}2^{-k}C_{k+1}}$ (21)
and
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{2k}\frac{{C_{k}}}{{C_{2k}}}}=3\binom{{2n}}{n}^{-1}\sum_{k=0}^{n}{\frac{{2^{k}\left({k+1}\right)}}{{\left({2\left(n-k\right)-1}\right)\left({2\left(n-k\right)-3}\right)}}\binom{{2\left({n-k}\right)}}{{n-k}}C_{k+1}}.$ (22)
Proof.
To derive (21), set $r=1$ and $s=3/2$ in (3) and use
$\binom{{2k+1}}{{3/2}}=\frac{4}{{3\sqrt{\pi}}}\frac{{\Gamma\left({2k+2}\right)}}{{\Gamma\left({2k+1/2}\right)}}=\frac{4}{{3\pi}}\frac{{2^{4k}}}{{C_{2k}}}$ (23)
and
$\binom{{n}}{{k-1/2}}=\frac{{\Gamma\left({n+1}\right)}}{{\Gamma\left({k+\frac{1}{2}}\right)\Gamma\left({n-k+\frac{3}{2}}\right)}}=\frac{{2^{2n+1}}}{{\pi\left({2\left({n-k}\right)+1}\right)}}\binom{{n}}{k}\binom{{2\left({n-k}\right)}}{{n-k}}^{-1}\binom{{2k}}{k}^{-1}.$ (24)
Identity (22) is obtained by setting r=1/2r=-1/2 and s=3/2s=-3/2 in (3) and using
$\binom{{2k-1/2}}{{-3/2}}=-2^{-4k-1}C_{2k}$ (25)
and
$\binom{{n-3/2}}{{k+1}}=\frac{{2^{-2k}}}{2}\frac{{\left({2\left({n-k}\right)-1}\right)\left({2\left({n-k}\right)-3}\right)}}{{\left({2n-1}\right)\left({k+1}\right)}}\binom{{2n}}{n}\binom{{n}}{k}\binom{{2\left({n-k}\right)}}{{n-k}}$. (26)
Proposition 2.
If $n$ is a non-negative integer and $r$ is a complex number that is not a non-positive integer, then
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}\frac{{2^{-2k}}}{{2k+r}}\,C_{k}}=\frac{{1}}{{n+1}}\binom{{n+r}}{{n+1}}^{-1}\sum_{k=0}^{n}{2^{-k}C_{k+1}\binom{{k+r-1}}{k}}$. (27)
In particular,
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}\frac{{2^{-2k}}}{{2k+1}}}\,C_{k}$ $=\frac{1}{{n+1}}\sum_{k=0}^{n}{2^{-k}C_{k+1}},$ (28)
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}\frac{{2^{-2k}}}{{4k+1}}}\,C_{k}$ $=\frac{{2^{2n}}}{{2n+1}}\binom{{2n}}{n}^{-1}\sum_{k=0}^{n}{2^{-3k}\binom{{2k}}{k}C_{k+1}},$ (29)
and
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}\frac{{2^{-2k}}}{{4k-1}}\,C_{k}}=2^{2n}\binom{{2n}}{n}^{-1}\left({2\sum_{k=1}^{n}{2^{-3k}C_{k-1}C_{k+1}}-1}\right).$ (30)
Proof.
Set $s=1$ in (3) and use
$\binom{{n+r-1}}{{k+r-1}}\binom{{k+r-1}}{k}=\binom{{n}}{k}\binom{{n+r-1}}{n}$. (31)
Identity (28) corresponds to setting $r=1$ in (27) while Identity (29) is obtained by evaluating (27) at $r=1/2$ and making use of
$\binom{{n+1/2}}{{n+1}}=\frac{{2n+1}}{{2^{2n+1}}}C_{n}$ (32)
and
$\binom{{k-1/2}}{k}=2^{-2k}\binom{{2k}}{k}$. (33)
Identity (30) is derived by shifting the summation index on the right-hand side of (27), setting r=1/2r=-1/2 and using
$\binom{{k-1/2}}{{k+1}}=-\frac{{C_{k}}}{{2^{2k+1}}}.$ (34)
Proposition 3.
If $n$ is a non-negative integer and $s$ is a complex number that is not a negative integer, then
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}}2^{-2k}C_{k}\binom{{2k+s}}{s}^{-1}=\frac{s}{{n+s}}\sum_{k=0}^{n}{\binom{{n}}{k}2^{-k}C_{k+1}\binom{{n-1+s}}{k}^{-1}}.$ (35)
In particular,
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}}2^{2k}C_{k}\binom{{4k}}{{2k}}^{-1}$ (36)
$\qquad=\frac{2^{n}3n}{{\left({n+1}\right)\left({n+2}\right)}}-\sum_{k=0}^{n-2}{\frac{{2^{k}}}{{2n-2k-1}}\binom{{n}}{k}^{2}C_{k+1}\binom{2k}{k}^{-1}\binom{2n}{2k}^{-1}}.$ (37)
Proof.
Set $r=s$ in (3). To derive (37), set s=1/2s=-1/2 in (35), write
$\sum_{k=0}^{n}{\binom{{n}}{k}2^{-k}C_{k+1}\binom{{n-2+1/2}}{k}^{-1}}$ $=\sum_{k=0}^{n-2}{\binom{{n}}{k}2^{-k}C_{k+1}\binom{{n-2+1/2}}{k}^{-1}}$ (38)
$\qquad+\binom{{n}}{{n-1}}2^{-(n-1)}C_{n}\binom{{n-2+1/2}}{{n-1}}^{-1}$ (39)
$\qquad\qquad+2^{-n}C_{n+1}\binom{{n-2+1/2}}{n}^{-1}$ (40)
and use
$\binom{{n-2+1/2}}{k}=2^{-2k}\binom{{2n-3}}{{2k}}\binom{{2k}}{k}\binom{{n-2}}{k}^{-1}$ (41)
and
$\binom{{2k-1/2}}{{-1/2}}=\binom{{2k-1/2}}{{2k}}=2^{-4k}\binom{{4k}}{{2k}}$. (42)
Proposition 4.
If $n$ and $r$ are non-negative integers and $s$ is a complex number such that sns-n is not a negative integer, then
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}\binom{{n-s-1}}{{2k+r}}^{-1}}=(-1)^{r-1}\,\frac{{n-s}}{{r+s}}\,\sum_{k=0}^{n}{\binom{{n}}{k}2^{-k}C_{k+1}\binom{{r+s-1}}{{k+r}}^{-1}}.$ (43)
In particular,
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\frac{{2^{2k}}}{{\left({2n-4k-1}\right)\left({2n-4k-3}\right)}}\binom{{n}}{{2k}}^{2}\binom{{2k}}{k}\binom{{2n}}{{4k}}^{-1}\frac{1}{{C_{2k}}}}=(-1)^{n}\,\frac{{2n+1}}{3}.$ (44)
Proof.
Write rnr-n for $r$ and sns-n for $s$ in (3) and write $r+s$ for $r$ in the resulting equation, noting that
$\binom{2k+r+s-n}{2k+r}=(-1)^{r}\binom{n-s-1}{2k+r}$. (45)
To derive (44), set $r=2$ and s=1/2s=-1/2 in (43) and use
$\binom{{1/2}}{{k+2}}=\frac{{(-1)^{k+1}}}{{2^{2k+3}}}C_{k+1}$ (46)
and
$\binom{{n-1/2}}{{2k+2}}=\frac{{2^{-4k}}}{8}\frac{{\left({2n-4k-1}\right)\left({2n-4k-3}\right)}}{{k+1}}\binom{{2n}}{{4k}}\binom{{n}}{{2k}}^{-1}C_{2k}.$ (47)
Proposition 5.
If $n$ is a non-negative integer and $r$ and $s$ are complex numbers that are not negative integers, then
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}H_{2k+r-s}\binom{{2k+r}}{s}^{-1}}$ (48)
$\qquad=\frac{s}{{n+r}}\sum_{k=0}^{n}{\binom{{n}}{k}2^{-k}C_{k+1}\left({H_{k+r-s}+H_{s-1}-H_{n+s-k-1}}\right)\binom{{n+r-1}}{{k+r-s}}^{-1}}.$ (49)
In particular,
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}H_{2k+r}}=2^{-n}C_{n+1}H_{n+r}-\sum_{k=0}^{n-1}{\binom{{n}}{k}\frac{2^{-k}}{n-k}C_{k+1}\binom{{n+r}}{{k+r}}^{-1}}.$ (50)
Proof.
Differentiate (3) with respect to $s$ to obtain
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}\left({H_{2k+r-s}-H_{s}}\right)\binom{{2k+r}}{s}^{-1}}$ (51)
$\qquad=-\frac{1}{{n+r}}\sum_{k=0}^{n}{\binom{{n}}{k}2^{-k}C_{k+1}\binom{{n+r-1}}{{k+r-s}}^{-1}}$ (52)
$\qquad\qquad+\frac{s}{{n+r}}\sum_{k=0}^{n}{\binom{{n}}{k}2^{-k}C_{k+1}\left({H_{k+r-s}-H_{n+s-k-1}}\right)\binom{{n+r-1}}{{k+r-s}}^{-1}},$ (53)
from which (49) follows after using (3) again. To derive (50), note that $sH_{s-1}=sH_{s}-1$, write the right-hand side of (49) as
$\frac{1}{{n+r}}\sum_{k=0}^{n-1}{\binom{{n}}{k}2^{-k}C_{k+1}\binom{{n+r-1}}{{k+r-s}}^{-1}\left({sH_{k+r-s}+sH_{s}-1-sH_{n+s-k-1}}\right)}$ (54)
$\qquad+\frac{1}{{n+r}}2^{-n}sC_{n+1}H_{n+r-s}\binom{{n+r-1}}{{n+r-s}}^{-1}$ (55)
and take the limit as $s$ approaches zero, using (4). ∎
Proposition 6.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}H_{2k}}=2^{-n}C_{n+1}H_{n}-\sum_{k=0}^{n-1}{\frac{{2^{-k}C_{k+1}}}{{n-k}}}$ (56)
and
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}O_{2k}}$ (57)
$\qquad=2^{-n}C_{n+1}O_{n}-2^{-n-1}\sum_{k=1}^{n}{\frac{{2^{3k}}}{k}\binom{{n}}{k}^{2}\binom{{2k}}{k}^{-1}\binom{{2n}}{{2k}}^{-1}C_{n-k+1}}.$ (58)
Proof.
Set $r=0$ in (50) to obtain (56), and r=1/2r=-1/2 to derive (58) after some algebra. ∎
Proposition 7.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-6k}C_{k}C_{2k}O_{2k}}$ (59)
$\qquad=\frac{1}{{2^{2n}\left({n+1}\right)}}\sum_{k=0}^{n}{\left({2\left({n-k}\right)+1}\right)\binom{{2\left({n-k}\right)}}{{n-k}}\binom{{2k}}{k}2^{-k}C_{k+1}\left({O_{k}-O_{n-k+1}+1}\right)}.$ (60)
Proof.
Set $r=1$ and $s=3/2$ in (49) and equate rational parts after using (23) and (24). ∎
3 A generalization involving Stirling numbers of the second kind
Lemma 2.
If $m$ and $u$ are non-negative integers, then
[…]
Proof.
Since
$\left({1-e^{x}}\right)^{u}=\sum_{p=0}^{u}{(-1)^{p}\binom{{u}}{p}e^{px}},$ (62)
we have
$\frac{{d^{m}}}{{dx^{m}}}{\left({1-e^{x}}\right)^{u}}=\sum_{p=0}^{u}{(-1)^{p}\binom{{u}}{p}p^{m}e^{px}};$ (63)
and hence (61) on account of (6).
Theorem 2.
If $n$ and $r$ are non-negative integers, then
[…]
Proof.
By writing $1/x$ for $x$, Identity (2) can also be written as
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}x^{n-2k}}=\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}2^{-k}C_{k+1}\left({1-x}\right)^{n-k}}.$ (65)
Now replace $x$ with $\exp x$, differentiate the resulting expression $r$ times with respect to $x$ and evaluate at $x=0$, making use of Lemma 2. ∎
4 Related results via binomial transform
Two sequences of complex numbers $(s_{k})$ and $(\sigma_{k})$, $k=0,1,2,\ldots$, are called a binomial- transform pair if
$s_{n}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\sigma_{k}},$ (66)
for every non-negative integer $n$. It is known that
$s_{n}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}\sigma_{k}}\iff\sigma_{n}=\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}s_{k}}.$ (67)
Lemma 3 ([1, Lemma 2.1]).
Let $\left\{(a_{k}),(\alpha_{k})\right\}$, $k=0,1,2,\ldots$, be a binomial-transform pair. Let $\mathcal{L}_{x}$ be a linear operator defined by $\mathcal{L}_{x}(x^{j})=a_{j}$ for every complex number $x$ and every non-negative integer $j$. Then $\mathcal{L}_{x}((1-x)^{j})=\alpha_{j}$.
Theorem 3.
Let $\left\{(a_{k}),(\alpha_{k})\right\}$, $k=0,1,2,\ldots$, be a binomial-transform pair. Then
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}2^{-k}C_{k+1}a_{n-k}}=\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}\alpha_{n-2k}}.$ (68)
Proof.
Write (65) as
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}2^{-k}C_{k+1}x^{n-k}}=\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}\left({1-x}\right)^{n-2k}}$ (69)
and apply Lemma 3. ∎
Remark 1.
It is sometimes useful to write (68) in the following form:
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{n-2k}C_{k}a_{n-2k}}$ (70)
$\qquad=\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{2k}C_{n-2k+1}\alpha_{2k}}-\sum_{k=1}^{\left\lceil{n/2}\right\rceil}{\binom{{n}}{{2k-1}}2^{2k-1}C_{n-2k+2}\alpha_{2k-1}}.$ (71)
Proposition 8.
If $n$ is a non-negative integer, then
[…]
Proof.
Use the binomial-transform pair $\{a_{k}=1,\alpha_{k}=\delta_{k0}\}$ in (68). Here $\delta_{ij}$ is Kronecker’s delta symbol. ∎
Remark 2.
Identity (72) was obtained by Suleiman and Sury [9].
Proposition 9.
If $n$ is a non-negative integer, then
[…]
and
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}\frac{{2^{n-2k}}}{{n-k+1}}\,C_{k+1}H_{n-k}}$ (74)
$\qquad=\begin{cases}-2\sum\limits_{k=0}^{(n-1)/2}{\dbinom{{n}}{{2k}}\dfrac{{2^{-2k}}}{{n-2k+1}}\,C_{k}O_{(n-2k+1)/2}},&\text{if $n$ is odd;}\\ 0,&\text{if $n$ is even.}\end{cases}$ (75)
Proof.
The second identity in [1, Proposition 13.3] is
[…]
where the binomial-transform pair $(a_{k})$ and $(\alpha_{k})$, $k=0,1,2,\ldots$, with
$a_{k}=\frac{2^{k}}{k+1}\text{ and }\alpha_{k}=\frac{1+(-1)^{k}}{2\,\left(k+1\right)},$ (77)
can be identified. Use these in (68) to obtain (73). From [1, Proposition 13.4]:
[…]
we recognize the binomial-transform pair $(a_{k})$ and $(\alpha_{k})$, $k=0,1,2,\ldots$, where
$a_{k}=\frac{2^{k}H_{k}}{k+1}\text{ and }\alpha_{k}=-\frac{1-(-1)^{k}}{\,\left(k+1\right)}\,O_{(k+1)/2}.$ (79)
Use of these in (68) gives (75). ∎
Proposition 10.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}\frac{{2^{2\left({n-2k}\right)}}}{{n-2k+1}}\,C_{k}}=\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}\frac{{2^{2k}}}{{2k+1}}\,C_{n-2k+1}}$ (80)
and
$\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}\frac{{2^{2\left({n-2k}\right)}}}{{n-2k+1}}\,C_{k}H_{n-2k}}=\frac{1}{2}\sum_{k=0}^{\left\lceil{n/2}\right\rceil}{\binom{{n}}{{2k-1}}\frac{{2^{2k}}}{k}\,C_{n-2k+2}O_{k}}.$ (81)
Proof.
Use (77) in (71) to obtain (80) and (78) to get (81). ∎
5 Additional results
The methods described in this paper can also be used to derive generalizations of another well-known combinatorial identity, Knuth’s old sum or the Reed Dawson identity:
[…]
Identity (82) occurs at $x=0$ in the following polynomial identity [2, Equation (6)]:
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}2^{-k}\binom{{2k}}{k}\left({1-x}\right)^{n-k}}=\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}\binom{{2k}}{k}x^{n-2k}}.$ (83)
Starting with the following variation on (83):
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-k}\binom{{2k}}{k}x^{k}}=\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}\binom{{2k}}{k}x^{2k}\left({1-x}\right)^{n-2k}},$ (84)
and proceeding as in the proof of Theorem 1 yields the following generalization of Knuth’s old sum.
Theorem 4.
If $r$ and $s$ are complex numbers that are not negative integers, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-k}\binom{{2k}}{k}\binom{{k+r}}{s}^{-1}}=\frac{s}{{n+r}}\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}\binom{{2k}}{k}\binom{{n+r-1}}{{2k+r-s}}^{-1}}.$ (85)
Equation (85) reduces to Knuth’s old sum (82) at $s=0$, while at $r=1=s$ we obtain the following binomial-transform identity:
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-k}C_{k}}=2^{-n}\binom{{n}}{{\left\lfloor{n/2}\right\rfloor}}$, (86)
after using the fact that
$\sum_{r=0}^{m}{2^{-2r}\binom{{2r}}{r}}=\sum_{r=0}^{m}{\binom{{r-1/2}}{r}}=\binom{{m+1/2}}{m}=\left({2m+1}\right)2^{-2m}\binom{{2m}}{m}$. (87)
Remark 3.
Identity (86) was derived on the MathWorld website [5].
Choosing $r=0$ and $s=1/2$ in (85) and using
$\binom{k}{1/2}=\frac{2^{2k+1}}{\pi}\,\binom{2k}{k}^{-1}$ (88)
and
$\binom{{n-1}}{{2k-1/2}}=\frac{{2^{2n}}}{{n\pi}}\binom{{n}}{{2k}}\binom{{4k}}{{2k}}^{-1}\binom{{2\left({n-2k}\right)}}{{n-2k}}^{-1},$ (89)
gives our next result.
Proposition 11.
If $n$ is a non-negative integer, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-3k}\binom{{2k}}{k}^{2}}=\frac{1}{{2^{2n}}}\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{2^{-2k}\binom{{2k}}{k}\binom{{4k}}{{2k}}\binom{{2\left({n-2k}\right)}}{{n-2k}}}$. (90)
The following variation on Equation (83):
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}2^{-k}\binom{{2k}}{k}x^{n-k}}=\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}\binom{{2k}}{k}\left(1-x\right)^{n-2k}},$ (91)
and the application of Lemma 3 leads to another generalization of Knuth’s old sum.
Theorem 5.
If $n$ and $r$ are non-negative integers, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-k}\binom{{2k}}{k}\left({n-k}\right)^{r}}$ (92)
[…]
Next, using (91) and Lemma 3, we present a related result to Knuth’s old sum involving a binomial-transform pair.
Theorem 6.
Let $\left\{(a_{k}),(\alpha_{k})\right\}$, $k=0,1,2,\ldots$, be a binomial-transform pair. Then
$\sum_{k=0}^{n}{(-1)^{n-k}\binom{{n}}{k}2^{-k}\binom{{2k}}{k}a_{n-k}}=\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}\binom{{2k}}{k}\alpha_{n-2k}}.$ (94)
Finally, by considering the following variation on identity (69):
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-k}C_{k+1}x^{k}}=\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}x^{2k}\left({1-x}\right)^{n-2k}},$ (95)
and working as in the proof of Theorem 1, we derive the next combinatorial identity.
Theorem 7.
If $n$ is a non-negative integer and $r$ and $s$ are complex numbers that are not negative integers, then
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-k}C_{k+1}\binom{{k+r}}{s}^{-1}}=\frac{s}{{n+r}}\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\binom{{n}}{{2k}}2^{-2k}C_{k}\binom{{n+r-1}}{{2k+r-s}}}^{-1}.$ (96)
In particular, setting $r=2$, $s=3/2$ and using the fact that
$\binom{k+2}{3/2}=\frac{16}{3\pi}\frac{2^{2k}}{C_{k+1}}$ (97)
and
$\binom{{n+1}}{{2k+1/2}}=\frac{{2^{2n+2}}}{\pi}\frac{{n+1}}{{\left({2n-4k+1}\right)\left({4k+1}\right)}}\binom{{n}}{{2k}}\binom{{4k}}{{2k}}^{-1}\binom{{2\left({n-2k}\right)}}{{n-2k}}^{-1},$ (98)
we obtain
$\sum_{k=0}^{n}{(-1)^{k}\binom{{n}}{k}2^{-3k}C_{k+1}^{2}}$ (99)
$\qquad=\frac{1}{{\left({n+1}\right)\left({n+2}\right)2^{2n-1}}}\sum_{k=0}^{\left\lfloor{n/2}\right\rfloor}{\left({2n-4k+1}\right)\left({4k+1}\right)2^{-2k}C_{k}\binom{{4k}}{{2k}}\binom{{2\left({n-2k}\right)}}{{n-2k}}}$. (100)